# A Single-Loop First-Order Algorithm for Linearly Constrained Bilevel Optimization

### v1

@@ 1 Introduction
In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter, ; franceschi2018bilevel, ; mackay2019self, ), data hypercleaning (shaban2019truncated, ), meta learning (rajeswaran2019meta, ; ji2020convergence, ), reinforcement learning (sutton2018reinforcement, ; hong2023two, ) and neural architecture search (liu2018darts, ; liang2019darts+, ); see survey papers (zhang2024introduction, ; liu2021investigating, ; sinha2017review, ) for additional discussions. While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation, ; liu2020generic, ; ji2021bilevel, ; dagreou2022framework, ; hong2023two, ; kwon2023fully, ), studies focusing on constrained BLO problems are relatively limited.
@@
For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$. This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized, ), hyperparameter optimization of constrained learning problems (xu2023efficient, ) and adversarial training (zhang2022revisiting, ) and draw significant attentions (tsaknakis2022implicit, ; khanduri2023linearly, ; kornowski2024first, ).
A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation, ; ji2021bilevel, ; ji2022will, ; chen2022single, ; khanduri2021near, ). For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ). However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ). The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
Recently, some first-order methods (kwon2023penalty, ; yao2024constrained, ; yao2024overcoming, ; jiang2024primal, ; kornowski2024first, ) have been proposed for addressing constrained BLO problems. Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem. For example, yao2024constrained ; yao2024overcoming reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem. However, neither yao2024constrained nor yao2024overcoming establishes clear relationships between the stationary points of their approximated problems and the original one.
@@
where $\Phi_{\delta}(x,y,z)=f(x,y)+\frac{1}{\delta}[g(x,y)-g(x,z)]$. Based on this reformulation, kwon2023penalty proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and jiang2024primal proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. However, kwon2023penalty only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration. Algorithms in jiang2024primal require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation. Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty, ; jiang2024primal, ) for coupled constraints $\mathcal{Y}(x)$.
@@
  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem. and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies). The simple single-loop structure also makes our algorithm easier to implement in practice compared to jiang2024primal .
@@
Table 1: Comparison of our paper with (kwon2023penalty, ; jiang2024primal, ; kornowski2024first, ). More detailed introductions and discussions of other related works can be found in Section˜2. Here, the “Complexity” for (kwon2023penalty, ; jiang2024primal, ) and SFLCB means the complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$; for (kornowski2024first, ), it means the complexity needed to achieve the $(\delta,\epsilon)$-Goldstein stationarity defined in kornowski2024first .
@@
kornowski2024first $Bx+Ay-b\leq 0$, LICQ holds $\tilde{O}(d\delta^{-1}\epsilon^{-3})$ double
kornowski2024first $Bx+Ay-b\leq 0$, LICQ holds $\tilde{O}(\delta^{-1}\epsilon^{-4})$ double
@@ 2 Related works
BLO without constraints. One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter, ). It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation, ). Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel, ; ji2022will, ; chen2022single, ; khanduri2021near, ; yang2023accelerating, ). Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient, ; grazzi2020iteration, ; liu2021towards, ; bolte2022automatic, ). Recently, penalty-based methods have gained traction as a promising approach for solving BLO. Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value, ; mehra2021penalty, ; liu2022bome, ; kwon2023fully, ; shen2023penalty, ; gao2022value, ; lu2024slm, ).
@@
BLO with constraints. There are two primary types of methods for solving constrained bilevel optimization problems. One is based on the implicit gradient method. Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly, ). tsaknakis2022implicit proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm. khanduri2023linearly proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$. xu2023efficient used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$. However, they only provided an asymptotic convergence analysis of their algorithm. xiao2023alternating proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem. However, these implicit gradient-type algorithms (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; xiao2023alternating, ) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
@@
Another commonly used approach for solving constrained BLO problems is based on penalty reformulation. For example, yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function. However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly. yao2024overcoming reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm. Compared to yao2024constrained , yao2024overcoming did not need the projection operator to the coupled constraint set. However, both yao2024constrained and yao2024overcoming did not establish very clear relationships between the stationary points of their approximated problems and the original one. For example, yao2024overcoming only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one. Recently, tsaknakis2023implicit ; jiang2024barrier proposed algorithms based on barrier approximation approach for constrained BLO problems. However, their algorithms also require the computations for the Hessian matrix.
@@ 4 Reformulation
Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit, ; khanduri2023linearly, ; xu2023efficient, ; jiang2024barrier, ; kwon2023penalty, ) to ensure the existence of $\nabla\Phi(x)$.
@@ 5 The SFLCB Algorithm
In the last section, we have justified the validity of our reformulation for coupled constrained BLO. In this section, we focus on a special and important case where the LL constraints are h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b. This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (yang2022decentralized, ; khanduri2023linearly, ), adversarial training (zhang2022revisiting, ; khanduri2023linearly, ), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6). For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (kwon2023penalty, ; jiang2024primal, ).
@@ 5.1 Convergence results
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA has full row rank, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply projected gradient descent (PGD) for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, we can find $\hat{v}$, $\hat{z}$ such that $\|\hat{v}-v^{*}(x_{0})\|\leq\delta$ and $\|\hat{z}-z^{*}(x_{0})\|\leq\delta$ with a complexity of $O(\log(\epsilon^{-1}))$. Set $y_{0}=z_{0}=\hat{z}$, $u_{0}=v_{0}=\hat{v}$, $\alpha_{0}=h(x_{0},y_{0})$, $\beta_{0}=h(x_{0},z_{0})$. With $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ as initial points and applying Algorithm˜1, we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$. Thus, the total complexity is $O(\epsilon^{-3}+\log(\epsilon^{-1}))=O(\epsilon^{-3})$.
@@ 6.2 Hyperparameter optimization in SVM
Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (xu2023efficient, ; yao2024constrained, ; yao2024overcoming, ; jiang2024primal, ). Here we consider the same problem formulation as in jiang2024primal , which formulates this problem as a coupled linearly constrained BLO problem. We conduct experiments comparing our SFLCB algorithm with GAM (xu2023efficient, ), LV-HBA (yao2024constrained, ), BLOCC (jiang2024primal, ), and BiC-GAFFA (yao2024overcoming, ) on the diabetes dataset (dua2017uci, ). Results are plotted in Figure˜2. We notice that our SFLCB algorithm converges significantly faster than other algorithms, which demonstrates the practical efficiency of the proposed SFLCB algorithm.
@@ 6.3 Transportation network design
We further conduct experiments on a transportation network design problem, following the same setting as in jiang2024primal . In this setting, we act as the operator, whose profit serves as the upper-level objective and is influenced by passenger behavior, which is modeled in the lower-level problem. Detailed formulations and settings can be found in Appendix˜F. We consider the two synthetic networks of 3 and 9 nodes, same as those considered in jiang2024primal . We compare SFLCB with BLOCC (jiang2024primal, ). Results are plotted in Figure˜3, which indicate that SFLCB significantly outperforms BLOCC on this network design task.
@@ 7 Conclusions and future directions
In this paper, for coupled constrained BLO problem in Equation˜1, we theoretically analyzed the relationship between the original hyper-objective $\Phi$ and the reformulated function $\Phi_{\delta}$ in Equation˜3, providing a solid justification for the validity of the reformulation. Especially, for the linearly constrained case, we proposed SFLCB, a single-loop, Hessian-free algorithm, improving the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ over previous works kwon2023penalty ; jiang2024primal . Our experiments on hyperparameter optimization for SVM and the transportation network design problem validated the practical efficiency of the proposed SFLCB algorithm. One limitation of our work is that the analysis is restricted to deterministic and linearly constrained settings. A promising direction for future research is to extend the current results to stochastic environments or more general constraint structures. Moreover, since the best-known complexity for first-order methods in unconstrained BLO (huang2024optimal, ) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.

### v2

@@ 1 Introduction
In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter; franceschi2018bilevel; mackay2019self), data hypercleaning (shaban2019truncated), meta learning (rajeswaran2019meta; ji2020convergence), reinforcement learning (sutton2018reinforcement; hong2023two) and neural architecture search (liu2018darts; liang2019darts+); see survey papers (zhang2024introduction; liu2021investigating; sinha2017review) for additional discussions. While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation; liu2020generic; ji2021bilevel; dagreou2022framework; hong2023two; kwon2023fully), studies focusing on constrained BLO problems are relatively limited.
@@
For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$. This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized), hyperparameter optimization of constrained learning problems (xu2023efficient) and adversarial training (zhang2022revisiting) and draw significant attentions (tsaknakis2022implicit; khanduri2023linearly; kornowski2024first).
A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation; ji2021bilevel; ji2022will; chen2022single; khanduri2021near). For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating). However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating). The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
Recently, some first-order methods (kwon2023penalty; yao2024constrained; yao2024overcoming; jiang2024primal; kornowski2024first) have been proposed for addressing constrained BLO problems. Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem. For example, yao2024constrained; yao2024overcoming reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem. However, neither yao2024constrained nor yao2024overcoming establishes clear relationships between the stationary points of their approximated problems and the original one.
@@
where $\Phi_{\delta}(x,y,z)=f(x,y)+\frac{1}{\delta}[g(x,y)-g(x,z)]$. Based on this reformulation, kwon2023penalty proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and jiang2024primal proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. However, kwon2023penalty only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration. Algorithms in jiang2024primal require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation. Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty; jiang2024primal) for coupled constraints $\mathcal{Y}(x)$.
@@
  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem, and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies). The simple single-loop structure also makes our algorithm easier to implement in practice compared to jiang2024primal.
@@
Table 1: Comparison of our paper with (kwon2023penalty; jiang2024primal). More detailed introductions and discussions of other related works can be found in Section˜2. Here, the “Complexity” means the iteration complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$ (3).
@@ 2 Related works
BLO without constraints. One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter). It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation). Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel; ji2022will; chen2022single; khanduri2021near; yang2023accelerating). Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient; grazzi2020iteration; liu2021towards; bolte2022automatic). Recently, penalty-based methods have gained traction as a promising approach for solving BLO. Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value; mehra2021penalty; liu2022bome; kwon2023fully; shen2023penalty; gao2022value; lu2024slm; lu2024first).
@@
BLO with constraints. There are two primary types of methods for solving constrained bilevel optimization problems. One is based on the implicit gradient method. Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly). tsaknakis2022implicit proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm. khanduri2023linearly proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$. xu2023efficient used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$. However, they only provided an asymptotic convergence analysis of their algorithm. xiao2023alternating proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem. However, these implicit gradient-type algorithms (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
@@
Another commonly used approach for solving constrained BLO problems is based on penalty reformulation. For example, lu2024first reformulated unconstrained and constrained BLO problems as structured minimax problems and introduced first-order methods with guarantees for finding $\epsilon$-KKT solutions. yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function. However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly. yao2024overcoming reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm. Compared to yao2024constrained, yao2024overcoming did not need the projection operator to the coupled constraint set. However, both yao2024constrained and yao2024overcoming did not establish very clear relationships between the stationary points of their approximated problems and the original one. For example, yao2024overcoming only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one. Recently, tsaknakis2023implicit; jiang2024barrier proposed algorithms based on barrier approximation approach for constrained BLO problems. However, their algorithms also require the computations for the Hessian matrix.
@@ 4 Reformulation
Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; jiang2024barrier; kwon2023penalty) to ensure the existence of $\nabla\Phi(x)$.
@@ 5 The SFLCB Algorithm
In the last section, we have justified the validity of our reformulation for coupled constrained BLO. In this section, we focus on a special and important case where the LL constraints are h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b. This particular category of constrained BLO problems encompasses a broad range of applications, including distributed optimization (yang2022decentralized; khanduri2023linearly), adversarial training (zhang2022revisiting; khanduri2023linearly), and hyperparameter optimization for constrained learning tasks such as hyperparameter optimization in SVM (see Section˜6). For this special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, we introduce a novel single-loop, first-order algorithm SFLCB, which achieves an improvement in the convergence rate compared to prior works (kwon2023penalty; jiang2024primal).
@@ 5.1 Convergence results
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA has full row rank, ˜3.1, 3.2, 3.3 hold, and $\delta=\Theta(\epsilon)\leq\mu_{g}/(2l_{f,1})$, if we apply projected gradient descent (PGD) for $\max_{v\in\mathbb{R}^{d_{h}}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},z)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, we can find $\hat{v}$, $\hat{z}$ such that $\|\hat{v}-v^{*}(x_{0})\|\leq\delta$ and $\|\hat{z}-z^{*}(x_{0})\|\leq\delta$ with a complexity of $O(\log(\epsilon^{-1}))$. Set $y_{0}=z_{0}=\hat{z}$, $u_{0}=v_{0}=\hat{v}$, $\alpha_{0}=h(x_{0},y_{0})$, $\beta_{0}=h(x_{0},z_{0})$. With $x_{0},y^{\prime}_{0},z^{\prime}_{0},u_{0},v_{0}$ as initial points and applying Algorithm˜1, we can find an $\epsilon$-stationary point of $\Phi_{\delta}$ with a complexity of $O(\epsilon^{-3})$. Thus, the total complexity is $O(\epsilon^{-3}+\log(\epsilon^{-1}))=O(\epsilon^{-3})$.
@@ 6.2 Hyperparameter optimization in SVM
Hyperparameter optimization in SVM is a well-known real-world application for constrained BLO problems that has been used in many prior works (xu2023efficient; yao2024constrained; yao2024overcoming; jiang2024primal). Here we consider the same problem formulation as in jiang2024primal, which formulates this problem as a coupled linearly constrained BLO problem. We conduct experiments comparing our SFLCB algorithm with GAM (xu2023efficient), LV-HBA (yao2024constrained), BLOCC (jiang2024primal), and BiC-GAFFA (yao2024overcoming) on the diabetes dataset (dua2017uci). Results are plotted in Figure˜2. We notice that our SFLCB algorithm converges significantly faster than other algorithms, which demonstrates the practical efficiency of the proposed SFLCB algorithm.
@@ 6.3 Transportation network design
We further conduct experiments on a transportation network design problem, following the same setting as in jiang2024primal. In this setting, we act as the operator, whose profit serves as the upper-level objective and is influenced by passenger behavior, which is modeled in the lower-level problem. Detailed formulations and settings can be found in Appendix˜F. We consider the two synthetic networks of 3 and 9 nodes, same as those considered in jiang2024primal. We compare SFLCB with BLOCC (jiang2024primal). Results are plotted in Figure˜3, which indicate that SFLCB significantly outperforms BLOCC on this network design task.
@@ 7 Conclusions and future directions
In this paper, for coupled constrained BLO problem in Equation˜1, we theoretically analyzed the relationship between the original hyper-objective $\Phi$ and the reformulated function $\Phi_{\delta}$ in Equation˜3, providing a solid justification for the validity of the reformulation. Especially, for the linearly constrained case, we proposed SFLCB, a single-loop, Hessian-free algorithm, improving the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ over previous works kwon2023penalty; jiang2024primal. Our experiments on hyperparameter optimization for SVM and the transportation network design problem validated the practical efficiency of the proposed SFLCB algorithm. One limitation of our work is that the analysis is restricted to deterministic and linearly constrained settings. A promising direction for future research is to extend the current results to stochastic environments or more general constraint structures. Moreover, since the best-known complexity for first-order methods in unconstrained BLO (huang2024optimal) is $O(\epsilon^{-2})$, it is also an interesting problem whether we can achieve this optimal rate in the constrained cases.